Find the volume of the solid formed by revolving the curve, r=a(1+cosθ), about the initial line.


1
Expert's answer
2021-06-09T10:32:01-0400

"V=\int\frac{2}{3} \pi*r^2 sin A" da

"a=\theta"

"\frac{2\pi}{3}∫a^3*(1-cos a)^3 sinA" da

"t=1-cos A"

Sin A "da=dt"

"a=t=0,"

"a=\pi,"

"t=2"

"Volume=\frac{8}{3} *\pi*a^3"